# Boosting drug-disease association prediction for drug repositioning via dual-feature extraction and cross-dual-domain decoding

### v3

@@ 10 DB01262 Decitabine CTD
TABLE VI: The top 10 candidate drugs for PD are ranked based on the DFDRNN prediction score
@@ 10 DB00215 Citalopram CTD
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table VI presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
IV Conclusion
@@
Acknowledgments
This work was supported by the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China under grants 62172072 and 62272115.

### v4

@@ 10 DB01262 Decitabine CTD
Table 6: The top 10 candidate drugs for PD are ranked based on the DFDRNN prediction score
@@ 10 DB00215 Citalopram CTD
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
4 Conclusion
@@
{acknowledgement}
The authors thank the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China (Grants 62172072 and 62272115) for their financial support.
5 Data Availability
The implementation and data of the DFDRNN model are available at https://github.com/siekersun/DFDRNN.git.
{suppinfo}
  • manuscript.pdf: Manuscript PDF File
  • manuscript.zip: This zip package contains files for an academic paper written in LaTeX, including the main document, references, and document class files for formatting.
  • supplementary.docx:Supporting Information for Publication